# Continuous transition and gapless roton inside fractional quantum anomalous Hall states

### v6

@@ Abstract
Roton is also a key elementary excitations in correlated quantum liquids Park et al. (2024), including the Bose–Einstein condensates (BECs) Landau (1941); Feynman and Cohen (1956), Fermi liquids Godfrin et al. (2012), and the FQH states (magnetoroton) Girvin et al. (1986); Kukushkin et al. (2009). FQAH states are believed to be both similar to and beyond the FQH physics in LLs, and recently, their collective excitations including the magnetoroton mode have been intensively discussed in lattice models and moiré materials Lu et al. (2024b, c); Long et al. (2025); Paul et al. (2025); Shen et al. (2025a); Kousa et al. (2025). Based on the knowledge of integer Hall crystals Tešanović et al. (1989) and FQH nematics Pu et al. (2024), we here propose a generic scheme for realizing FQAH+CDW state as illustrated in Fig. 1. Starting from a symmetric FQAH state, if the magenetoroton gap is softened and closed at some finite momentum by tuning interactions while keeping the charge gap finite, the ground state becomes an FQAH+CDW state with the same Hall conductance and spontaneously generated moderate CDW order. Stronger CDW order will eventually lead to topologically trivial states Tešanović et al. (1989); Pu et al. (2024). As the roton mode is introduced in the superfluid liquid helium and the FQH states are often interpreted as the superfluid of composite bosons Zhang et al. (1989), our scenario is analogous to supersolids triggered by the rotons in superfluids Tanzi et al. (2019); Böttcher et al. (2019); Chomaz et al. (2019); Zhang et al. (2019); Alaña et al. (2023). Although the intuitive picture is transparent, to the best of our knowledge, no clear realization in a concrete microscopic model for either FQH or FQAH states has been reported, except for the topologically trivial CDWs triggered by rotons Kumar and Bhatt (2022); Lu et al. (2024c); Shen et al. (2025a).
@@
$H=$ $-\sum_{\langle i,j\rangle}te^{i\phi_{ij}}(c_{i}^{\dagger}c^{\ }_{j}+h.c.)-\sum_{\langle\hskip-1.42262pt\langle i,j\rangle\hskip-1.42262pt\rangle}t^{\prime}_{ij}(c_{i}^{\dagger}c^{\ }_{j}+h.c.)$ (1)
$-\sum_{\langle\hskip-1.42262pt\langle\hskip-1.42262pt\langle i,j\rangle\hskip-1.42262pt\rangle\hskip-1.42262pt\rangle}t^{\prime\prime}(c_{i}^{\dagger}c^{\ }_{j}+h.c.)+\sum_{i,j}V_{ij}n_{i}n_{j},$
@@
with : $t=1$ (as the energy unit), $t^{\prime}_{ij}=\pm 1/(2+\sqrt{2})$, $t^{\prime\prime}=-1/(2+2\sqrt{2})$ and $\phi_{ij}=\frac{\pi}{4}$, as shown in Fig. 2 (a). The two bands have non-trivial Chern number $C=\pm 1$ and the lower Chern band is almost flat with these optimal hopping parameters Sheng et al. (2011); Sun et al. (2011). We further consider the nearest-neighbor (NN) interaction $V_{1}$, the next-nearest-neighbor interaction $V_{2}$, and the third nearest-neighbor interaction $V_{3}$. The results in the main text are based on exact diagonalization (ED) and infinite density matrix renormalization group (iDMRG) McCulloch (2008). For the iDMRG simulations, we mainly consider the $N_{y}\times N_{x}\times 2=3\times 2\times 2$-site unit cells (i.e. 6-leg infinite cylinder) and we keep up to $D=3000$ $U(1)$ states. The maximum truncation error is $\sim 10^{-9}$ with good convergence.
@@
Figure 3: Continuous FQAH-FQAHS transition. iDMRG results at different bond dimensions of a $N_{y}=3$ cylinder with infinite length (a) Bipartite entanglement entropy and (b) correlation length as functions of $V_{3}$, which both diverge at the critical point. The inset of panel (a) shows the fitting of central charge at critical point $V_{3}^{c}=0.61$ according to $S_{\mathrm{E}}=\frac{c}{6}\log(\xi)$ and the fitted result is $c\sim 0.7$. The power-law divergence of correlation length approaching the critical point at the (c) FQAH and (d) FQAHS sides are shown. When the bond dimension increases, the scalings of correlation length are both approaching $|V_{3}-V_{3}^{c}|^{-\nu}$ with $\nu=1$ of the 2D Ising universality class. (e) Smectic CDW order parameter and (f) structure factor at $(\pi,0)$ as functions of $V_{3}$. The inset of panel (e) in the double-logarithm scale shows the fitting of the order parameter according to $\delta_{\mathrm{smectic}}\sim(V_{3}-V_{3}^{c})^{\beta}$ and the fitted exponent is $\beta\sim 0.14$. $S(\pi,0)$ that measures the smectic fluctuations also tends to diverge at the critical point and quickly drops after the establishment of smectic order.
@@
We focus on the $\nu=2/3$ filling of the flat Chern band. The phase diagram of considering only $V_{3}$ interaction has been reported Lu et al. (2024d). The ground state is a Fermi liquid at the non-interacting limit, and is a PSM at the large-$V_{3}$ limit. The PSM is a non-Fermi liquid with anisotropic transport properties and irrelevant inter-stripe tunnelings Emery et al. (2000); Lu et al. (2024d); Li et al. (2024). Interestingly, between the Fermi liquid and the PSM states, there is an FQAHS (FQAH + smectic CDW) state. Although the ground-state and thermodynamic properties of the FQAHS state has been well understood, its origin and the possible connection to any symmetric FQAH state remain unexplored. Besides, in another work at the same $\nu=2/3$ filling, the $V_{1}-V_{2}$ phase diagram has been studied Lu et al. (2024e). There is a symmetric FQAH state with competing interactions, surrounded by different CDW states. Although the global $V_{1}-V_{2}-V_{3}$ phase diagram has not been studied, the discoveries above provide an opportunity to explore the connections between the symmetric FQAH state and the FQAHS state with coexisting orders.
@@
Continuous FQAH-FQAHS transition.— The bipartite entanglement entropy ($S_{\mathrm{E}}$) and correlation length ($\xi$) around the first transition are shown in Fig. 3 (a,b), respectively. They both diverge with increasing bond dimensions at the critical point $V_{3}^{c}=0.61$, supporting the continuous nature of the transition. Since the FQAHS state with the same topological order as the FQAH state breaks additional discrete translation symmetry and has a $(\pi,0)$ smectic CDW order and doubled ground-state degeneracy, we expect that it belongs to the Ising universality class. In the inset of Fig.3(a), we extrapolate the central charge at the critical point according to $S_{\mathrm{E}}=\frac{c}{6}\log(\xi)$. The extrapolated $c\sim 0.7$ is roughly consistent with the expected $c=0.5$ for the 2D Ising transition (due to the quasi-1D geometry of infinite cylinder in iDMRG simulations). We further show the behavior of correlation length as a function of $|V_{3}-V_{3}^{c}|$ when approaching the critical point from both sides in Fig. 3 (c,d) respectively. When the bond dimension of the iDMRG simulations increases, the correlation length scales closer to the power law $\xi\sim|V_{3}-V_{3}^{c}|^{-\nu}$ and the fitted exponent is approaching to $\nu=1$, which further confirms the continuous transition and supports its Ising nature. As illustrated in Ref. Lu et al. (2024d), the $(\pi,0)$ smectic order exists in only one sublattice (B sublattice here) in the FQAHS state. Therefore, we define the smectic CDW order parameter as $\delta_{\mathrm{smectic}}=\frac{2}{N^{\prime}}\sum_{i}^{{}^{\prime}}(-1)^{x_{i}}n^{\mathrm{B}}_{\mathbf{r_{i}}}$ with the $\sum^{\prime}$ summation over $N^{\prime}(=N_{x}N_{y})$ sites. Besides, the static structure factor of density-density correlation functions is defined as $S(\mathbf{q})=\sum_{j}e^{-i\mathbf{q}(\mathbf{r_{j}}-\mathbf{r_{0}})}(\langle n_{j}^{\mathrm{B}}n_{0}^{\mathrm{B}}\rangle-\langle n_{j}^{\mathrm{B}}\rangle\langle n_{0}^{\mathrm{B}}\rangle)$, with $S(\pi,0)$ as the smectic fluctuations. The smectic CDW order and the smectic fluctuations as functions of $V_{3}$ are shown in Fig. 3 (e,f), respectively. When approaching the critical point from the FQAH state, the smectic fluctuations gradually increase and peak at $V_{3}^{c}$ which will eventually diverge with bond dimensions. After the spontaneously established smectic order at the critical point, the smectic fluctuations decrease in the FQAHS state. In the inset of Fig. 3 (e), we show the smectic order (with bond dimension $D=3000$) as a function of $V_{3}-V_{3}^{c}$ in the double-logarithm scale. We find that the power-law scaling of the order parameter $\delta_{\mathrm{smectic}}\sim(V_{3}-V_{3}^{c})^{\beta}$ with the fitted exponent $\beta\sim 0.14$, which is very close to the expected exponent $\beta=\frac{1}{8}$ for the 2D Ising transition.
@@
Considering the fast-developing cold-atom, dipolar-gas and quantum-circuit platforms in realising FQ(A)H states Yao et al. (2013); Grass et al. (2018); L茅onard et al. (2023); Wang et al. (2024); Shen et al. (2025b), we argue that our work provides a generic mechanism and a further strong movitation to explore the anisotropic transport and coexisting CDW orders in such states. We notice that the spectroscopic scheme based on two interfering Laguerre-Gaussian beams to detect the magnetoroton and edge excitations in such systems has been proposed Binanti et al. (2024), and the roton-driven topologically trivial crystals have been reported for such as interacting bosons in the gauge fields Mukherjee et al. (2022). Although the example in our work is by tuning longer-range interactions in a fermionic FQAH state on the lattice, the generic mechanism by softening the roton mode is similar in bosonic and continuous systems, and the similar techniques in the BECs such as the periodic potentials favoring the target CDW momenta could be used. Besides, we have shown that this roton-driven FQ(A)H-FQ(A)H+CDW transition can be continuous, which is beneficial to the experimental evolutions compared with the first-order transitions (which could lead to the generation of high-energy excitations upon dynamically crossing the crystallization) as in the superfluid-supersolid evolutions Zhang et al. (2019); Alaña et al. (2023).
@@
Moreover, our scenario shares another similarity that the interactions finally lead to topologically trivial CDW states with stronger order parameter. While the topological transition from the integer Hall crystal to Wigner crystal is understood from the further closing single-particle gap Tešanović et al. (1989), the transitions from the FQAHS state here, or the FQH nematics, to topologically trivial and gapless CDW states (PSM here) has not been fully understood Pu et al. (2024). In the previous work with only $V_{3}$ interaction, the ground-state FQAHS-PSM transition is first-order, but the finite-temperature one is possibly continuous Lu et al. (2024d). Therefore, it is worth investigating whether there exists a continuous FQAHS-PSM transition at zero temperature, which could be a potentially continuous topological transition beyond the traditional Ginzburg-Landau paradigm. We leave this interesting possiblity for future work.
A further interesting question is what protects the robust charge gap when the neural gap goes soft. In the original proposal of integer Hall crystal, although the tuning interaction is two-body, the four-body interaction is necessary to stablize the charge gap in that case Tešanović et al. (1989). But in our FQAH-FQAHS transition as well as the FQH-nematic FQH transition Yang (2020); Pu et al. (2024), two-body interactions are sufficient. A deeper answer is still elusive. Besides, how to soften the graviton mode in FQAH states is also an interesting open question.
@@ Acknowledgements.
Acknowledgments —  We thank Kai Sun and Bo Yang for helpful discussions. We thank Eduardo Fradkin about the Ising nature of this transition and Zhendong Zhang for discussions on ultracold systems. We use TeNPy package for the infinite DMRG simulations Hauschild and Pollmann (2018). HYL and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. AoE/P-701/20, 17309822, HKU C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22). We thank HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong, as well as the Beijng PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported within this paper. H.Q. Wu acknowledge the support from National Natural Science Foundation of China (No. 12474248), GuangDong Basic and Applied Basic Research Foundation (No. 2023B1515120013), Guangdong Fundamental Research Center for Magnetoelectric Physics (Grant No.2024B0303390001), Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (No. 2022B1212010008).
Supplementary Information for:
Continuous transition and gapless roton inside fractional quantum anomalous Hall states
I Supplementary Information for
Continuous transition and gapless roton inside fractional quantum anomalous Hall states
II A. ED spectra of fixed $V_{1}=1.1$ and $V_{2}=1$
Figure S1: Energy spectra with fixed $V_{1}=1.1$ and $V_{2}=1$. These energy levels are obtained from ED simulations of $3\times 4\times 2$ tori. The red dashed lines are the phase boundaries as in the Fig. 2 (b) of the main text. The FQAH state has a 3-fold ground-state degeneracy and the FQAHS state has a 6-fold degeneracy, while the PSM state is gapless. The supporting spectra of each phase under twisted boundary conditions are also shown in Fig. 2 (b). The first phase boundary is from the iDMRG results in the main text, while the second phase boundary is determined from the level crossing at $V_{3}\sim 2.84$.
Figure S2: ED results of the FQAH-FQAHS transition on a different path. (a) and (b) show the Brilliou zone of $3\times 4\times 2$ torus and $3\times 6\times 2$ torus respectively. The momentum sectors related to the ground states are marked with light/dark green/blue colors, while the others are marked with gray. (c) and (d) show the energy spectra with the changing of $\lambda$ for $3\times 4\times 2$ torus and $3\times 6\times 2$ torus respectively, and the colors represent the momentum sectors specified in (a) and (b). In the FQAHS phase, the number of fully degenerate states in periodic boundary conditions are labeled. (e) and (f) show the twisted energy spectrum at $\lambda=0.2$ and $\lambda=0.4$, respectively, exhibiting the 3-fold (6-fold) ground-state degeneracy of FQAH(S) state. The charge gap with the changing of $\lambda$ for a $3\times 4\times 2$ torus is shown in (g). (h) The second derivative of ground state energy with respective to $\lambda$.
In this section, we show the supplementary energy spectra of ED simulations in the same parameter path (fixed $V_{1}=1.1$ and $V_{2}=1$) as in the main text.
Although the spectra of each phase under twisted boundary conditions are shown in Fig. 2 (b), we show the evolution of spectra (without twisting the boundary conditions) when tuning $V_{3}$ in Fig.S1. The first phase boundary is from the iDMRG results in the main text, and it is also consistent with the ED spectra. Although it is under periodic boundary conditions, it is still clear that the neutral gap closes and re-opens around $V_{3}\sim 0.61$, and we will explain more about the evolution of ground states in the next section. The second phase boundary is determined from the level crossing at $V_{3}\sim 2.84$. Note that, in our ED simulations, the repulsive interactions are written as $V_{ij}\sum_{ij}(n_{i}-1/2)(n_{j}-1/2)$, which leads to some simultaneous shift of the exact values but not the relative values of the energies.
III B. The FQAH-FQAHS transition on a different path
[…]
The charge gap defined as $\Delta_{\mathrm{charge}}=E(N,N_{\mathrm{e}}+1)+E(N,N_{\mathrm{e}}-1)-2E(N,N_{\mathrm{e}})$ in a $3\times 4\times 2$ torus along the same path of changing $\lambda$ is shown in Fig. S2 (g), which is consistent with the iDMRG results in the main text that although the neutral excitation gap is closed at the critical point, the charge excitation remains gapped. The second derivatives of the ground state energy (taken from the state with lowest energy) with respective to the tuning parameter $\lambda$ for the two system sizes are shown in Fig. S2 (h), which also supports that this FQAH-FQAHS transition is continuous, which can also be seen from the smooth merging of the ground states around the transition point in Fig. S2 (c) and (d).

### v7

@@ Abstract
Roton is also a key elementary excitations in correlated quantum liquids Park et al. (2024), including the Bose–Einstein condensates (BECs) Landau (1941); Feynman and Cohen (1956), Fermi liquids Godfrin et al. (2012), and the FQH states (magnetoroton) Girvin et al. (1986); Kukushkin et al. (2009). FQAH states are believed to be both similar to and beyond the FQH physics in LLs, and recently, their collective excitations including the magnetoroton mode have been intensively discussed in lattice models and moiré materials Lu et al. (2024b, c); Long et al. (2025); Paul et al. (2025); Shen et al. (2025a); Kousa et al. (2025). Based on the knowledge of integer Hall crystals Tešanović et al. (1989) and FQH nematics Pu et al. (2024), we here propose a generic scheme for realizing FQAH+CDW state as illustrated in Fig. 1. Starting from a symmetric FQAH state, if the magnetoroton gap is softened and closed at some finite momentum by tuning interactions while keeping the charge gap finite, the ground state becomes an FQAH+CDW state with the same Hall conductance and spontaneously generated moderate CDW order. Unlike the integer Hall crystal and nematic FQH states in Landau levels, the neutral excitation of the FQAH+CDW state here is gapped since we consider discrete translational symmetry. Stronger CDW order will eventually lead to topologically trivial states Tešanović et al. (1989); Pu et al. (2024). As the roton mode is introduced in the superfluid liquid helium and the FQH states are often interpreted as the superfluid of composite bosons Zhang et al. (1989), our scenario is analogous to supersolids triggered by the rotons in superfluids Tanzi et al. (2019); Böttcher et al. (2019); Chomaz et al. (2019); Zhang et al. (2019); Alaña et al. (2023). Although the intuitive picture is transparent, to the best of our knowledge, no clear realization in a concrete microscopic model for either FQH or FQAH states has been reported, except for the topologically trivial CDWs triggered by rotons Kumar and Bhatt (2022); Lu et al. (2024c); Shen et al. (2025a).
@@
$H=$ $-\sum_{\langle i,j\rangle}te^{i\phi_{ij}}(c_{i}^{\dagger}c^{\ }_{j}+h.c.)-\sum_{\langle\thinspace\langle i,j\rangle\thinspace\rangle}t^{\prime}_{ij}(c_{i}^{\dagger}c^{\ }_{j}+h.c.)$ (1)
$-\sum_{\langle\thinspace\langle\thinspace\langle i,j\rangle\thinspace\rangle\thinspace\rangle}t^{\prime\prime}(c_{i}^{\dagger}c^{\ }_{j}+h.c.)+\sum_{i,j}V_{ij}n_{i}n_{j},$
@@
with : $t=1$ (as the energy unit), $t^{\prime}_{ij}=\pm 1/(2+\sqrt{2})$, $t^{\prime\prime}=-1/(2+2\sqrt{2})$ and $\phi_{ij}=\frac{\pi}{4}$, as shown in Fig. 2 (a). The two bands have nontrivial Chern number $C=\pm 1$ and the lower Chern band is almost flat with these optimal hopping parameters Sheng et al. (2011); Sun et al. (2011). We further consider the nearest-neighbor (NN) interaction $V_{1}$, the next-nearest-neighbor interaction $V_{2}$, and the third nearest-neighbor interaction $V_{3}$. The results in the main text are based on exact diagonalization (ED) and infinite density matrix renormalization group (iDMRG) McCulloch (2008). For the iDMRG simulations, we mainly consider the $N_{y}\times N_{x}\times 2=3\times 2\times 2$-site unit cells (i.e. 6-leg infinite cylinder) and we keep $D=3000$ $U(1)$ states for most simulations. The maximum truncation error is $\sim 10^{-9}$ with good convergence.
@@
Figure 3: Continuous FQAH-FQAHS transition. iDMRG results at different bond dimensions of a $N_{y}=3$ cylinder with infinite length (a) Bipartite entanglement entropy and (b) correlation length as functions of $V_{3}$, which both diverge at the critical point. The inset of panel (a) shows the fitting of central charge at critical point $V_{3}^{c}=0.61$ according to $S_{\mathrm{E}}=\frac{c}{6}\log(\xi)$ and the fitted result from the data points of largest bond dimensions is $c\sim 0.6$ (we simulated up to $D=4000$ for this part). The power-law divergence of correlation length approaching the critical point at the (c) FQAH and (d) FQAHS sides are shown. When the bond dimension increases, the scalings of correlation length are both approaching $|V_{3}-V_{3}^{c}|^{-\nu}$ with $\nu=1$ of the 2D Ising universality class. (e) Smectic CDW order parameter and (f) structure factor at $(\pi,0)$ as functions of $V_{3}$. The inset of panel (e) in the double-logarithm scale shows the fitting of the order parameter according to $\delta_{\mathrm{smectic}}\sim(V_{3}-V_{3}^{c})^{\beta}$ and the fitted exponent is $\beta\sim 0.14$. $S(\pi,0)$ that measures the smectic fluctuations also tends to diverge at the critical point and quickly drops after the establishment of smectic order.
@@
We focus on the $\nu=2/3$ filling of the flat Chern band. The phase diagram of considering only $V_{3}$ interaction has been reported Lu et al. (2024d). The ground state is a Fermi liquid at the non-interacting limit, and is a PSM at the large-$V_{3}$ limit. The PSM is a non-Fermi liquid with anisotropic transport properties and irrelevant inter-stripe tunnelings Emery et al. (2000); Lu et al. (2024d); Li et al. (2024). Interestingly, between the Fermi liquid and the PSM states, there is an FQAHS (FQAH + smectic CDW) state. Although the ground-state and thermodynamic properties of the FQAHS state have been well understood, its origin and the possible connection to any symmetric FQAH state remain unexplored. In another work at the same $\nu=2/3$ filling, the $V_{1}-V_{2}$ phase diagram has been studied Lu et al. (2024e). There is a symmetric FQAH state with competing interactions, surrounded by different CDW states. However, whether there exists a direct transition between the symmetric FQAH state and the FQAHS state with coexisting orders in the global phase diagram is still not clear.
@@
Continuous FQAH-FQAHS transition.— The bipartite entanglement entropy ($S_{\mathrm{E}}$) and correlation length ($\xi$) around the first transition are shown in Fig. 3 (a,b), respectively. They both diverge with increasing bond dimensions at the critical point $V_{3}^{c}=0.61$, supporting the continuous nature of the transition. Since the FQAHS state with the same topological order as the FQAH state breaks additional discrete translation symmetry and has a $(\pi,0)$ smectic CDW order and doubled ground-state degeneracy, we expect that it belongs to the Ising universality class. In the inset of Fig.3(a), we extrapolate the central charge at the critical point according to $S_{\mathrm{E}}=\frac{c}{6}\log(\xi)$ (we have considered up to $D=4000$ for this part). The extrapolated $c\sim 0.6$ is roughly consistent with the expected $c=0.5$ for the 2D Ising transition (due to the quasi-1D geometry of infinite cylinder in iDMRG simulations). We further show the behavior of correlation length as a function of $|V_{3}-V_{3}^{c}|$ when approaching the critical point from both sides in Fig. 3 (c,d) respectively. When the bond dimension of the iDMRG simulations increases, the correlation length scales closer to the power law $\xi\sim|V_{3}-V_{3}^{c}|^{-\nu}$ and the fitted exponent is approaching to $\nu=1$, which further confirms the continuous transition and supports its Ising nature. As illustrated in Ref. Lu et al. (2024d), the $(\pi,0)$ smectic order exists in only one sublattice (B sublattice here) in the FQAHS state. Therefore, we define the smectic CDW order parameter as $\delta_{\mathrm{smectic}}=\frac{2}{N^{\prime}}\sum_{i}^{{}^{\prime}}(-1)^{x_{i}}\langle n^{\mathrm{B}}_{\mathbf{r_{i}}}\rangle$ with the $\sum^{\prime}$ summation over $N^{\prime}(=N_{x}N_{y})$ sites. Besides, the static structure factor of density-density correlation functions is defined as $S(\mathbf{q})=\sum_{j}e^{-i\mathbf{q}(\mathbf{r_{j}}-\mathbf{r_{0}})}(\langle n_{j}^{\mathrm{B}}n_{0}^{\mathrm{B}}\rangle-\langle n_{j}^{\mathrm{B}}\rangle\langle n_{0}^{\mathrm{B}}\rangle)$, with $S(\pi,0)$ as the smectic fluctuations. The smectic CDW order and the smectic fluctuations as functions of $V_{3}$ are shown in Fig. 3 (e,f), respectively. When approaching the critical point from the FQAH state, the smectic fluctuations gradually increase and peak at $V_{3}^{c}$ which will eventually diverge with bond dimensions. After the spontaneously established smectic order at the critical point, the smectic fluctuations decrease in the FQAHS state due to the explicitly translation symmetry breaking with spatially variant electron distribution. In the inset of Fig. 3 (e), we show the smectic order (with bond dimension $D=3000$) as a function of $V_{3}-V_{3}^{c}$ in the double-logarithm scale. We find that the power-law scaling of the order parameter $\delta_{\mathrm{smectic}}\sim(V_{3}-V_{3}^{c})^{\beta}$ with the fitted exponent $\beta\sim 0.14$, which is very close to the expected exponent $\beta=\frac{1}{8}$ for the 2D Ising transition.
@@
Considering the fast-developing cold-atom, dipolar-gas and quantum-circuit platforms in realising FQ(A)H states Yao et al. (2013); Grass et al. (2018); Léonard et al. (2023); Wang et al. (2024); Shen et al. (2025b), we argue that our work provides a generic mechanism and a further strong movitation to explore the anisotropic transport and coexisting CDW orders in such states. We notice that the spectroscopic scheme based on two interfering Laguerre-Gaussian beams to detect the magnetoroton and edge excitations in such systems has been proposed Binanti et al. (2024), and the roton-driven topologically trivial crystals have been reported for such as interacting bosons in the gauge fields Mukherjee et al. (2022). Although the example in our work is by tuning longer-range interactions in a fermionic FQAH state on the lattice, the generic mechanism by softening the roton mode is similar in bosonic and continuous systems, and the similar techniques in the BECs such as the periodic potentials favoring the target CDW momenta could be used. Besides, we have shown that this roton-driven FQ(A)H-FQ(A)H+CDW transition can be continuous, which is beneficial to the experimental evolutions compared with the first-order transitions (which could lead to the generation of high-energy excitations upon dynamically crossing the crystallization) as in the superfluid-supersolid evolutions Zhang et al. (2019); Alaña et al. (2023).
@@
Moreover, our scenario shares another similarity that the interactions finally lead to topologically trivial CDW states with stronger order parameter. While the topological transition from the integer Hall crystal to Wigner crystal is understood from the further closing single-particle gap Tešanović et al. (1989), the transitions from the FQAHS state here, or the FQH nematics, to topologically trivial and gapless CDW states (PSM here) has not been fully understood Pu et al. (2024). In the previous work with only $V_{3}$ interaction, the ground-state FQAHS-PSM transition is first-order, but the finite-temperature one is possibly continuous Lu et al. (2024d). Therefore, it is worth investigating whether there exists a continuous FQAHS-PSM transition at zero temperature, which could be a potentially continuous topological transition beyond the traditional Ginzburg-Landau paradigm. We leave this interesting possibility for future work.
A further interesting question is what protects the robust charge gap when the neural gap goes soft. In the original proposal of integer Hall crystal, although the tuning interaction is two-body, the four-body interaction is necessary to stabilize the charge gap in that case Tešanović et al. (1989). But in our FQAH-FQAHS transition as well as the FQH-nematic FQH transition Yang (2020); Pu et al. (2024), two-body interactions are sufficient. A deeper answer is still elusive. Besides, how to soften the graviton mode in FQAH states is also an interesting open question.
In this work, the studied systems do not have exact $C_{4}$ rotational symmetry due to the choice of geometries. We believe that, in the $C_{4}$ symmetric systems, the translation symmetry breaking along x and y directions should be degenerate and would further enlarge the ground-state degeneracy in the FQAHS phase. While such system sizes are beyond our current numerical capability, it is interesting for future explorations.
@@ Acknowledgements.
Acknowledgments —  We thank Kai Sun and Bo Yang for helpful discussions. We thank Eduardo Fradkin about the Ising nature of this transition and Zhendong Zhang for discussions on ultracold systems. We use TeNPy package for the infinite DMRG simulations Hauschild and Pollmann (2018). HYL and ZYM acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. AoE/P-701/20, 17309822, HKU C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22). We thank HPC2021 system under the Information Technology Services and the Blackbody HPC system at the Department of Physics, University of Hong Kong, as well as the Beijing Paratera Tech Corp., Ltd par for providing HPC resources that have contributed to the research results reported within this paper. H.Q. Wu acknowledge the support from National Natural Science Foundation of China (No. 12474248), GuangDong Basic and Applied Basic Research Foundation (No. 2023B1515120013), Guangdong Fundamental Research Center for Magnetoelectric Physics (Grant No.2024B0303390001), Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (No. 2022B1212010008).